OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = -1/2 and the point (-2, -3).

OpenStudy (anonymous):

y =mx + b -3 = -1/2 (-2) + b -3 = 1 + b -4 = b y = -1/2x - 4

OpenStudy (jdoe0001):

in standard form. <---

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

y+4 + 1/2 x = 0

OpenStudy (anonymous):

can you help me with one more

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

thanks. Input the equation of the given line in standard form. https://media.glynlyon.com/g_alg02_ccss_2013/3/ui3.gif

OpenStudy (anonymous):

y = -1/3x + 1 y + 1/3x = 1

OpenStudy (anonymous):

hey, sorry but for the first question, the answer is y + 1/2x = -4, not y + 4 + 1/2 x = 0